### v1

@@ 4.1 Offensive Language
is defined as a speech targeted at a specific group of people with the intention of causing harm, violence, or social disruption Waseem et al. (2017); Sigurbergsson and Derczynski (2020); ajvazi2022dataset, regardless of the language used. Common examples of hate speech include the use of racial and homophobic slurs. Waseem et al. established a list of rules to identify hate speech, including negative statements or stereotypes about minority groups, disparaging or racial terms, and support for entities that encourage hate speech. Hate speech is usually done based on characteristics such as race, color, ethnicity, gender, sexual orientation, nationality, or religion. For instance, in Chinese, the term ”婚驴” /the married donkey/ is commonly used, particularly by radical feminists to describe married women, which can be seen as a form of derogatory language targeting married women.
@@
Adult content, intended for 18 and older audience, encompasses various materials, from sexual content to violence and explicit language Mubarak et al. (2021). Not all adult content is inherently offensive; its offensiveness often depends on context and cultural norms. In the Chinese context, adult content is conveyed implicitly due to strict regulations. This has led to the widespread use of subversive phrases in offensive language zidani2018represented. For example, terms like “萝卜”(carrot) are used to reference male genitals, and “奶子”(milk) is employed to refer to female breasts. These metaphorical expressions pose several challenges to the automatic detection of adult content.
@@
Sarcasm and dark humor are two linguistic phenomena that have been studied extensively in linguistics Skalicky and Crossley (2018), psychology Kreuz and Glucksberg (1989), and NLP Joshi et al. (2018). While these forms of speech can serve as sources of entertainment in non-offensive contexts, they are more frequently encountered in potentially offensive situations. Sarcasm presents a controversial aspect of offensive language, largely due to its varying interpretations. In this paper, we adopt the definition from merriam-webster: ”a sharp and often satirical or ironic utterance designed to cut or give pain”. [^1] https://www.merriam-webster.com/dictionary/sarcasm Using this definition, we categorize sarcasm that fits this description as offensive language, given its intent to cause harm. While previous research hasn’t explicitly included sarcasm as a type of offensive language, several studies link sarcasm to negative sentiments like anger or jealousy Riloff et al. (2013); 2020.osact-1.5. Failing to classify sarcasm as offensive language can overlook its potential harm Frenda et al. (2022a, b). Hence, we include sarcasm as a form of offensiveness.
Conversely, dark humor entails employing humor to address taboo or morbid subjects, as noted in a study by Mittal et al. (2021) and said2023phenomenon. However, dark humor can cross into offensiveness when it turns cruel or targets marginalized or vulnerable groups, including those with disabilities, specific genders, races, sexualities, religions, or cultural identities kumari2023emoffmeme. Social media platforms frequently feature dark humor, with memes incorporating such humor being widely shared kiela2020hateful.
@@ 5.3 Toxic Content Dataset
In the realm of Chinese offensive language detection, the TOXICN dataset represents a significant milestone lu2023facilitating. This dataset was meticulously constructed by collecting data from prominent online platforms such as Zhihu and Tieba, followed by a series of filtering and labeling measures. [^4] https://www.zhihu.com/, http://c.tieba.baidu.com/ The resulting dataset features a sophisticated multi-level labeling system, encompassing categories like offensive language, hate speech, targeted groups, and expression types. This comprehensive approach offers researchers valuable insights into the complexities and nuances of offensive language in the Chinese context. TOXICN was used to introduce a benchmark known as ”Toxic Knowledge Enhancement (TKE)” designed to enhance the detection of toxic language using lexical knowledge. TKE involves enriching the representation of sentences by incorporating lexical features. This is achieved by embedding each token of a sentence in a vector space and introducing toxic knowledge through a specific embedding, which categorizes tokens as related to insults. The enhanced token representation combines the toxic embedding with the original word embedding, enabling a more effective integration of toxic and linguistic information. TKE can be applied to any pre-trained large language models to enhance the detection of toxic language.
@@ 6 Current Approaches and Models
These models are built not only using supervised Jiang et al. (2021) and semi-supervised Wu et al. (2019) approaches but also exploring the use of adversarial learning to solve this task Liu et al. (2020a). Most of the supervised approaches such as those presented by wu2019detecting and Yuan et al. (2019), provide a more adapted approach to identifying offensive language in a text. However, they too have their limitations. For instance, wu2019detecting’s use of BILSTM models may face challenges in handling long-range dependencies found in social media posts. Yuan et al. (2019)’s model trained on WeiboHate, can suffer from imbalanced class distribution and the variability of annotator judgments, potentially impacting the quality of the training data. Several other models were presented in the literature that made use of ML techniques for classifying Chinese text as offensive or not.
@@ 8.3 Neologism in Chinese
Neologism presents a unique challenge for detecting offensive language in Chinese. These are linguistic expressions, words, or phrases that are often used to circumvent censorship or to express taboo or sensitive topics in a covert manner Subversive. Even worse, people would embed implicit offense inside a seemingly polite and acceptable phrase. These phrases are often deeply embedded in specific cultural contexts and may be difficult to detect without an understanding of the relevant cultural references and language nuances. An example of subversive expressions from a post in Baidu Tieba, a popular Chinese online forum platform where users can discuss various topics and share information, is given in Figure LABEL:figure:tieba. [^8] https://tieba.baidu.com/

### v3

@@ 4.1 Offensive Language
is defined as a speech targeted at a specific group of people with the intention of causing harm, violence, or social disruption Waseem et al. (2017); Sigurbergsson and Derczynski (2020); Ajvazi and Hardmeier (2022), regardless of the language used. Common examples of hate speech include the use of racial and homophobic slurs. Waseem et al. established a list of rules to identify hate speech, including negative statements or stereotypes about minority groups, disparaging or racial terms, and support for entities that encourage hate speech. Hate speech is usually done based on characteristics such as race, color, ethnicity, gender, sexual orientation, nationality, or religion. For instance, in Chinese, the term ”婚驴” /the married donkey/ is commonly used, particularly by radical feminists to describe married women, which can be seen as a form of derogatory language targeting married women.
@@
Adult content, intended for 18 and older audience, encompasses various materials, from sexual content to violence and explicit language Mubarak et al. (2021). Not all adult content is inherently offensive; its offensiveness often depends on context and cultural norms. In the Chinese context, adult content is conveyed implicitly due to strict regulations. This has led to the widespread use of subversive phrases in offensive language Zidani (2018). For example, terms like “萝卜”(carrot) are used to reference male genitals, and “奶子”(milk) is employed to refer to female breasts. These metaphorical expressions pose several challenges to the automatic detection of adult content.
@@
Sarcasm and dark humor are two linguistic phenomena that have been studied extensively in linguistics Skalicky and Crossley (2018), psychology Kreuz and Glucksberg (1989), and NLP Joshi et al. (2018). While these forms of speech can serve as sources of entertainment in non-offensive contexts, they are more frequently encountered in potentially offensive situations. Sarcasm presents a controversial aspect of offensive language, largely due to its varying interpretations. In this paper, we adopt the definition from merriam-webster: ”a sharp and often satirical or ironic utterance designed to cut or give pain”. [^1] https://www.merriam-webster.com/dictionary/sarcasm Using this definition, we categorize sarcasm that fits this description as offensive language, given its intent to cause harm. While previous research hasn’t explicitly included sarcasm as a type of offensive language, several studies link sarcasm to negative sentiments like anger or jealousy Riloff et al. (2013); Abu Farha and Magdy (2020). Failing to classify sarcasm as offensive language can overlook its potential harm Frenda et al. (2022a, b). Hence, we include sarcasm as a form of offensiveness.
Conversely, dark humor entails employing humor to address taboo or morbid subjects, as noted in a study by Mittal et al. (2021) . However, dark humor can cross into offensiveness when it turns cruel or targets marginalized or vulnerable groups, including those with disabilities, specific genders, races, sexualities, religions, or cultural identities Kumari et al. (2023). Social media platforms frequently feature dark humor, with memes incorporating such humor being widely shared Kiela et al. (2021).
@@ 5.3 Toxic Content Dataset
In the realm of Chinese offensive language detection, the TOXICN dataset represents a significant milestone Lu et al. (2023). This dataset was meticulously constructed by collecting data from prominent online platforms such as Zhihu and Tieba, followed by a series of filtering and labeling measures. [^4] https://www.zhihu.com/, http://c.tieba.baidu.com/ The resulting dataset features a sophisticated multi-level labeling system, encompassing categories like offensive language, hate speech, targeted groups, and expression types. This comprehensive approach offers researchers valuable insights into the complexities and nuances of offensive language in the Chinese context. TOXICN was used to introduce a benchmark known as ”Toxic Knowledge Enhancement (TKE)” designed to enhance the detection of toxic language using lexical knowledge. TKE involves enriching the representation of sentences by incorporating lexical features. This is achieved by embedding each token of a sentence in a vector space and introducing toxic knowledge through a specific embedding, which categorizes tokens as related to insults. The enhanced token representation combines the toxic embedding with the original word embedding, enabling a more effective integration of toxic and linguistic information. TKE can be applied to any pre-trained large language models to enhance the detection of toxic language.
@@ 6 Current Approaches and Models
These models are built not only using supervised Jiang et al. (2021) and semi-supervised Wu et al. (2019) approaches but also exploring the use of adversarial learning to solve this task Liu et al. (2020a). Most of the supervised approaches such as those presented by Yuan et al. (2019), provide a more adapted approach to identifying offensive language in a text. However, they too have their limitations. For example,Yuan et al. (2019)’s model trained on WeiboHate, can suffer from imbalanced class distribution and the variability of annotator judgments, potentially impacting the quality of the training data. Several other models were presented in the literature that made use of ML techniques for classifying Chinese text as offensive or not.
@@ 8.3 Neologism in Chinese
Neologism presents a unique challenge for detecting offensive language in Chinese. These are linguistic expressions, words, or phrases that are often used to circumvent censorship or to express taboo or sensitive topics in a covert manner Wiener (2011). Even worse, people would embed implicit offense inside a seemingly polite and acceptable phrase. These phrases are often deeply embedded in specific cultural contexts and may be difficult to detect without an understanding of the relevant cultural references and language nuances. An example of subversive expressions from a post in Baidu Tieba, a popular Chinese online forum platform where users can discuss various topics and share information, is given in Figure LABEL:figure:tieba. [^8] https://tieba.baidu.com/